OpenStudy (anonymous):

tan inverse (cos x / 1-sin x) express in simplest form

OpenStudy (saifoo.khan):

Justin Bieber Sucks

OpenStudy (mimi_x3):

Saifoo~ sucks hehe

OpenStudy (saifoo.khan):

Mimi Sucks more than Jb. Even she copies her photo instead of Micheal Jackson

OpenStudy (mimi_x3):

http://www.wolframalpha.com/input/?i=tan+inverse+%28cos+x+%2F+1-sin+x%29 wtf , i did not copy , i didnt even know mans You stalker

OpenStudy (saifoo.khan):

Ever Green Link @eshaan

OpenStudy (mimi_x3):

Fail link hehe

OpenStudy (saifoo.khan):

See meme there^

OpenStudy (mimi_x3):

Gawd , i have to change my dp then argh Thanks to you :P

OpenStudy (anonymous):

Yup@Saif....

OpenStudy (mimi_x3):

I look better in it though xD

OpenStudy (saifoo.khan):

haha! that dosnt matter. secondly, why are you trying to get in place of MJ?? thirdly, looks dosn't matter much. :

OpenStudy (mimi_x3):

LOL, tbh , i didnt even know there was one with MJ xDD Well, the the app i used stole MJ's one not my fault hehe Gawd, im tired *sigh*

OpenStudy (anonymous):

hey can just help me out in this prob!!!

OpenStudy (saifoo.khan):

No, why you have jb on your photo?

OpenStudy (mimi_x3):

I just gave you the link

OpenStudy (mimi_x3):

Too~ sleepy to solve it for you hehe Ishaan can do it xD

OpenStudy (saifoo.khan):

GOD* yea, that could happen... the app must have stolen your photo.

OpenStudy (mimi_x3):

Yes, and im innocent xD

OpenStudy (saifoo.khan):

LOL, good joke.

OpenStudy (mimi_x3):

That wasn't even a joke hehe Such a fail

OpenStudy (mimi_x3):

Im still curious how did you search my pic ?

OpenStudy (anonymous):

\[\tan^{-1}\frac{sin(\frac{\pi}{2} -x)}{1- cos(\frac{\pi}{2}-x)}\] \[\frac{\pi}{2}-x = y \] \[\tan^{-1}\frac{siny}{1-cosy}\] \[\tan^{-1} \frac{2*sin\frac{y}{2}*cos\frac{y}{2}}{2sin^2\frac{y}{2}}\]

OpenStudy (anonymous):

Now you can do it ...

OpenStudy (saifoo.khan):

LOL, i never said that "IT IS A JOKE". MAFIA.

OpenStudy (mimi_x3):

How did you search my pic, STALKER

OpenStudy (saifoo.khan):

first, i got the info about your location. btw, have you even been to Bora Bora?

OpenStudy (saifoo.khan):

Haha, that is NOT stalking. Idk if you do it for others.

OpenStudy (anonymous):

Lol

OpenStudy (mimi_x3):

Info about my location ?! STALKER !?

OpenStudy (saifoo.khan):

HAHA, that is tracking not stalking. Dumb.

OpenStudy (saifoo.khan):

Mafia Tracking.

OpenStudy (mimi_x3):

Same thing Stupid Idiot argh I shouldnt have told you where i live xD

OpenStudy (saifoo.khan):

Haha, that wont matter either.. coz i already knew where you did. :P Please dont tell your things here.

OpenStudy (mimi_x3):

I didn't say much thoguh hehe I only said I live in Sydney Gawd, i regret telling you where i live STALKER

OpenStudy (anonymous):

hey will u plzzz help me out from this prob plz

OpenStudy (mimi_x3):

Ishaan just helped you :P

OpenStudy (saifoo.khan):

whatever, that dosn't even matter. LOL, that is not stalking, you are a astalker.

OpenStudy (saifoo.khan):

stalker*

OpenStudy (anonymous):

hey can u just complete it

OpenStudy (mimi_x3):

Well, I didn't stalk your pic or anything hehe How did you find it, im curious Tell me before I go to sleep please

OpenStudy (saifoo.khan):

Firstly, that is not called stalker. secondly, how's i know you did that or not. finally, you will get a message while you are sleeping.

OpenStudy (anonymous):

You can do it ....on your own now ..I am not going to give you answer you need to use your head ... and also because you are JB fan...

OpenStudy (mimi_x3):

LOL *hater alert* xD Tell me or i wont be able to sleep xD

OpenStudy (saifoo.khan):

haha @ i-shaan. lol STOP hating me, then it's great. :D

OpenStudy (mimi_x3):

Ofc i h8 , tell me pls , or im going to die of curiosity hehe

OpenStudy (saifoo.khan):

It's a Golden opportunity then! :D

OpenStudy (anonymous):

hello sir it is not the final answer right i want the final answer.

OpenStudy (mimi_x3):

LOL, look at the link I gave you Saifoo, please tell me *begs*

OpenStudy (saifoo.khan):

he is saying he will not answer more bcz you have a jb pic

OpenStudy (mimi_x3):

Just change your pic arjunan and ishaan wil help you

OpenStudy (anonymous):

hey the link gives me the derivative of that

OpenStudy (mimi_x3):

LOL, idk , I didn't even look at it. Just look at Ishaans and work it ouy